What is the length, to the nearest tenth of an inch, of the longest board that can fit through a rectangular doorway 6 feet and 8 inches in height and 3 feet in width?
Answer by rfer(16322)

6.66^2+3^2=c^2
44.36+9=53.36
sqrt 53.36=7.3 ft=7ft 3.6in